My daughter has this example : 2/3 x -5 =7
she has several problems like this to solve but missing one of the above pieces of information.

Answer by ankor@dixie-net.com(22740)

: (2/3)x - 5 = 7
:
It's a good idea to mult by a value that will get rid of the denominator
Mult thru by 3 and you have:
2x - 15 = 21
;
Add 15 to both sides to isolate the unknown on the left:
2x = 21 + 15
:
2x = 36
:
Divide both sides by 2
x = 36/2
x = 18
:
:
It's a good idea to check your solution by substituting the solution in the original equation:
(2/3)18 - 5 = 7
12 - 5 = 7
: